OpenStudy (anonymous):

The function y1(x) = x is a solution of x^2y''+xy'-y=0. Find the general solution of the differential equation x^2y''+xy'-y=2x.

OpenStudy (anonymous):

When I plugged in the equations into WA, I did not get that as a solution, so I am really confused on what to do.

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=x%5E2y%27%27%2Bxy%27-y%3D2x Interesting ... there seem to be complex term on the solution

OpenStudy (experimentx):

the constants of integration could be complex valued, just replace it by c_2, rest should be fine.

OpenStudy (anonymous):

there is an equation that this problem references but I do not understand how it applies.

OpenStudy (experimentx):

this is Euler Cauchy equation ... there is along article on wikipedia about how to solve it ... also since one solution is given, try using the method of reduction of order to solve it http://en.wikipedia.org/wiki/Reduction_of_order

OpenStudy (anonymous):

that example on there is kind of confusing.

OpenStudy (experimentx):

Let, \( y_2 = v(x) x \) be the other solution, not put that solution on your DE

OpenStudy (experimentx):

x^ (v(x) x) '' + ... = 2x

OpenStudy (anonymous):

is v(x) the homogeneous equation

OpenStudy (anonymous):

or i guess i dont understand what v(x) actually is

OpenStudy (experimentx):

no ... it's just assumption that your other solution is of that form.

OpenStudy (anonymous):

so how do you solve for the other solution when you dont even know what v(x) actually is?

OpenStudy (experimentx):

hold on for a while.

OpenStudy (anonymous):

It the problem says to reference another problem which states. Let y=y1(x) be a solution of (H):y''+p(x)y'+q(x)y=0 where p and q are continuous function on an interval I. Let a be in I and assume that y1(x) not equal to 0 on I. Set y2(x) = y1(x)integral (e^-(integral p(u)du))/y1^2(t)) dt. Show that y2 is a solution of (H) and that y1 and y2 are linearly independent. I see that it is the same kind of problem, but I dont see how this helps in anyway.

OpenStudy (experimentx):

you should be getting in this form http://upload.wikimedia.org/math/2/7/e/27ec56e8d0dc1f64f12feb692f4e5660.png which is a linear DE

OpenStudy (experimentx):

Of course substituting \( y_2 = v x \) gives \[x^2 (v'' x + 2 v' )+x(v'x+v) - vx = 2x \\ x^3 v'' +3x^2 v' = 2x \\ x^2 v'' + 3x v' = 2\] Let v' = u, we get \( x^2 u' + 3x u = 2\) which is linear ... now solve for 'u' then sovle for 'v' .. and put that thing in \(y = vx \) and get your final solution.

OpenStudy (anonymous):

am i just going to get u by itself and then integrate?

OpenStudy (experimentx):

yep ... we know v' = u, so we can find v from u.

OpenStudy (experimentx):

this is pretty hectic ... you should have directly done this which is much simpler http://en.wikipedia.org/wiki/Cauchy%E2%80%93Euler_equation

OpenStudy (anonymous):

so i integrate 1/3(2-x^2)x^-1 and get x^2/3 - x^4/12

OpenStudy (anonymous):

but im not sure where the u went?

OpenStudy (anonymous):

if u = (2-x^2(u'))/3x i guess i ignored the u, but you cant do that

OpenStudy (experimentx):